A 17-kg sled is being pulled along the horizontal snow-covered ground by a horizontal force of 22 N. Starting from rest, the sled attains a speed of 2.2 m/s in 8.3 m. Find the coefficient of kinetic friction between the runners of the sled and the snow.
Answer #1

here,

mass , m = 17 kg

horizontal force , F = 22 N

let the coefficient of friction be uk

accelration , a = ( F - uk*m*g)/m

a = ( 22 - uk *17*9.8)/17

a = 1.294 - 9.8*uk ....(1)

speed atained , v = 2.2 m/s

s = 8.3 m

using third equation of motion

v^2 - u^2 = 2*a*s

2.2^2 - 0 = 2 * ( 1.294 - 9.8 *uk) * 8.3

uk = 0.102

the coefficient of kinetic friction between the runners of the sled and the snow is 0.102